4. Suppose thatz~N(0,1) (a) Find the probability that z >1. (b) Find the probability that zS-1.96. (c) Find the probability that 0>>-2. (d) Find the value之c such that there is a 90% probability that-2cくZく
Answer #1

0-1587 1-0.8413 = 0.158 Mslo 0.1587. o. 9구 50 겨 -0.9チ50 U50の Ce) 〉,군 O 500-0228 = 0.47720-9 o-oi -근 を -·モe 0.0 5- ー1.6 45 えーー1.645 = 1.645 o 5 Ze-= 1.645